# A Hybrid Multi-Factor Network with Dynamic Sequence Modeling for Early Warning of Intraoperative Hypotension

### v2

@@ Abstract
Intraoperative hypotension (IOH) prediction using past physiological signals is crucial, as IOH can lead to inadequate organ perfusion, increasing the risk of severe complications and mortality. However, existing IOH prediction methods often rely on static modeling, overlooking the complex temporal dependencies and non-stationary nature of physiological signals. In this paper, we propose a Hybrid Multi-Factor (HMF) network that models IOH prediction as a dynamic sequence forecasting problem, explicitly capturing temporal dependencies and physiological non-stationarity.. Specifically, we formalize physiological signal dynamics as a sequence of multivariate time series, and decompose them into trend and seasonal components, enabling distinct modeling of long-term and periodic variations. For each component, we employ a patch-based Transformer encoder to extract representative features with the concern of computational efficiency and representation quality. Furthermore, to mitigate distributional drift arising from the evolving signals, we introduce a symmetric normalization mechanism. Extensive experiments on both a publicly available dataset and a private dataset collected from real-world hospital settings demonstrate that our approach significantly outperforms competitive baselines. We hope HMF offers a new perspective on IOH prediction and further enhances surgical safety [^2] https://github.com/Mingyue-Cheng/HMF.
@@ 1 Introduction
Figure 1: Illustration of MAP Sequence: Temporal Dynamics and Distribution Shifts in Intraoperative Hypotension Event.
@@
Intraoperative mortality has decreased by a factor of $100$ over the past century, making deaths during surgery a rare occurrence Li et al. (2009); Xue et al. (2022). However, mortality within the first month following surgery remains a significant concern, with approximately 2% of patients undergoing inpatient noncardiac surgery dying within $30$ days postoperatively Spence et al. (2019)—amounting to more than $4$ million deaths worldwide each year Saugel and Sessler (2021). These postoperative deaths are most strongly linked to complications, which are often triggered by intraoperative events. Among these, intraoperative hypotension (IOH) is a common and serious complication Kim et al. (2023), characterized by a significant drop in Mean Arterial Pressure (MAP) [^3] MAP, derived from Arterial Blood Pressure (ABP), reflects tissue perfusion and is calculated as $MAP=DBP+\frac{1}{3}(SBP-DBP)$, where SBP and DBP denote peak and minimum arterial pressures per cardiac cycle. A sustained decline in SBP and DBP reduces MAP, risking organ hypoperfusion. sustained over a period of time.
@@
Previous studies Hatib et al. (2018a) have shown that IOH events can be predicted using machine learning applied to physiological signals. Clinical trials further indicate that early warnings and timely interventions help mitigate hypotension severity and reduce postoperative complications Hwang et al. (2023); Fernandes et al. (2021). Recent advances in IOH prediction primarily follow a feature extraction-based approach Lee et al. (2021); Davies et al. (2020), where handcrafted features are used with classifiers like logistic regression and random forests. However, these methods face two key limitations: (1) reliance on handcrafted features limits their ability to capture temporal dependencies, and (2) their models lack the flexibility to adapt to complex real-world IOH scenarios. Additionally, IOH definitions vary across clinical settings, and fixed threshold-based criteria may fail to accurately characterize IOH, particularly in hypertensive or elderly patients, where more nuanced definitions are often needed.
@@
To overcome the limitations of existing approaches, we reformulate IOH prediction as a multivariate time series forecasting problem. The core rationale behind this formulation is that time series modeling effectively captures the temporal dynamics and distribution shifts inherent in physiological signals. Since hypotensive events are defined based on MAP series, accurate sequence forecasting directly impacts risk identification, offering greater flexibility. However, as illustrated in Figure 1, several challenges arise in MAP series prediction. First, the temporal dynamic waveform of MAP consists of multiple components, including trends and periodic patterns, which may need to be modeled separately to improve forecasting accuracy. Second, the high information redundancy caused by the extended length of MAP waveforms increases computational complexity and complicates representation learning in sequence modeling. Finally, evolving distribution shifts introduce additional challenges, as sudden fluctuations in blood pressure alter the statistical properties (e.g., mean and standard deviation) of the MAP series, making it difficult to maintain consistent predictive performance. Addressing these challenges is essential for developing a high-accurate and adaptable IOH prediction framework.
@@
To address these limitations, we introduce the Hybrid Multi-Factor (HMF) framework, which explicitly models temporal dependencies and mitigates distribution shifts, ensuring adaptive and accurate IOH prediction. A key feature of HMF is its use of sequence decomposition techniques to model the trend and seasonal components separately, allowing for a more structured representation of historical dynamics. By integrating multi-factor physiological signals, including MAP and SBP, HMF effectively captures both local fluctuations and long-term dependencies. To efficiently extract informative features, we adopt a patch-based Transformer encoder to learn representations of the given sequence. Additionally, a symmetric normalization mechanism is introduced to mitigate distribution shifts caused by evolving physiological signals. We conduct extensive experiments on both public and private datasets collected from real-world clinical scenarios. The results demonstrate the superiority of HMF in improving predictive performance while effectively modeling multi-factor temporal interactions.
The key contributions are summarized as follows:
@@
Figure 3: Proposed Hybrid Multi-Factor Model Architecture for Intraoperative Hypotension Prediction.
@@ 2 Preliminaries
To formulate the IOH prediction task, we first define how physiological features are extracted and structured, which serves as the foundation for subsequent problem formulation.
2.1 Multi-factor Feature Construction
To predict intraoperative hypotension (IOH), real-time monitoring and modeling techniques analyze arterial blood pressure (ABP) dynamics. Identifying early patterns preceding hypotensive events is crucial for timely intervention. This study structures multi-factor ABP-derived features to enhance forecasting accuracy.
During surgery, ABP is continuously monitored at high frequency (e.g., 100 Hz), capturing hemodynamic fluctuations. Prior studies reveal a strong correlation between ABP variations within a context window and subsequent hypotension. Clinically, hypotension is defined based on Mean Arterial Pressure (MAP), the average pressure per cardiac cycle, considered hypotensive when sustained below a threshold (e.g., one minute). Systolic Blood Pressure (SBP), the peak pressure per cycle, is included due to its physiological link to MAP. These ABP-derived signals—MAP and SBP—serve as core predictive features.
To construct informative representations, we extract continuous ABP segments, removing outliers and artifacts. A sliding window approach processes ABP series, extracting SBP (peak pressure) and Diastolic Blood Pressure (DBP, lowest pressure). MAP is computed as:
$\text{MAP}=\frac{1}{3}\times\text{SBP}+\frac{2}{3}\times\text{DBP}.$ (1)
Hypotensive events are identified when MAP remains below 65 mmHg for at least one minute, following clinical guidelines. Event annotation details, including threshold selection and temporal alignment, are further provided in the experimental and supplementary material.
For IOH prediction, we construct a multivariate time series dataset incorporating SBP and MAP to capture both short-term fluctuations and long-term trends. This formulation enables effective feature extraction and temporal alignment, as illustrated in Figure 2. By integrating multiple ABP-derived factors, our approach improves the model’s ability to learn temporal dependencies and physiological dynamics associated with IOH.
@@ 2.2 Problem Definition
Building on the multi-factor feature construction framework, we formulate IOH risk prediction as a dynamic sequence forecasting problem. Given real-time arterial blood pressure (ABP) monitoring, our goal is to predict future mean arterial pressure (MAP) trends, enabling early warnings for potential hypotensive events. Formally, let $\mathbf{X}_{\text{MAP}}=\{x_{\text{MAP}}^{1},x_{\text{MAP}}^{2},\dots,x_{\text{MAP}}^{T}\}$ and $\mathbf{X}_{\text{SBP}}=\{x_{\text{SBP}}^{1},x_{\text{SBP}}^{2},\dots,x_{\text{SBP}}^{T}\}$ denote the historical MAP and systolic blood pressure (SBP) sequences over a context window of length $T$. The objective is to forecast the future MAP values $\mathbf{\hat{Y}}_{\text{MAP}}=\{\hat{y}_{\text{MAP}}^{T+1},\dots,\hat{y}_{\text{MAP}}^{T+\tau}\}$ over a prediction horizon $\tau$. A hypotensive event occurs if the predicted MAP falls below a predefined threshold $\theta_{\text{MAP}}$. Further details on event annotation and threshold selection are provided in supplement material part. By leveraging multi-factor temporal dependencies, our formulation enhances the reliability of IOH risk assessment, facilitating timely clinical intervention.
@@ 3 The Proposed HMF
This section provides a detailed description of the proposed HMF framework.
@@ 3.2 Symmetric Normalization
where $\mu_{X}$ and $\sigma_{X}$ are the mean and standard deviation of $X$. This transformation allows the model to focus on relative changes rather than absolute values, reducing the impact of abrupt fluctuations and improving pattern recognition. Additionally, by normalizing across instances, the method mitigates inter-patient variability, ensuring stable feature distributions.
@@ 3.3 Sequence Decomposition Preprocessing
3.4 Dynamic Feature Representation Layer
@@
To model the intricate dynamics of MAP and SBP series during surgery, we apply a patch embedding technique after the decomposition layer. This reduces sequence length, lowering computational complexity and improving efficiency. Patch-based modeling captures local patterns while mitigating noise and outliers, enhancing prediction robustness.
@@
$W_{\text{patch}}=\text{Conv1D}_{3}(W),$ (6)
[…]
@@
Effective IOH prediction requires capturing both short-term and long-term dependencies in MAP and SBP series. The Transformer encoder, with its self-attention mechanism, is well-suited for this task, dynamically weighting the importance of different time segments to model complex temporal patterns.
@@ 3.5 Prediction layer
To ensure that the temporal patterns captured by the Transformer encoder are effectively utilized in MAP series predictions, we employ a linear layer to map the sequence representation $Z$ to the predicted output. This transformation enables the model to preserve sequence dependencies while efficiently translating extracted features into precise forecasts. To maintain temporal coherence, we employ a recurrent forecasting strategy, where each predicted step conditions the next, ensuring consistency across the forecasted sequence. The mapping from $Z\in\mathbb{R}^{\frac{L}{S}\times d_{\text{model}}}$ to $Y\in\mathbb{R}^{\frac{L}{S}\times\frac{2TS}{L}}$ is formulated as:
@@ 3.6 Optimization Strategies
To enhance prediction accuracy, we incorporate a patch-based autoregressive method. The process begins with the final patch of the decomposed component $W$, denoted as $Y_{0}=W_{\frac{L}{S}}$, and sequentially generates subsequent patches using:
@@ 4.1 Experimental Setup
We conduct experiments on two real-world datasets: CH-OPBP and VitalDB Lee (2018). The CH-OPBP dataset is collected from real-world intraoperative monitoring records of 3,422 patients, sampled at 100Hz. The data acquisition process strictly adheres to the ethical guidelines and protocols approved by the hospital ethics committee, ensuring compliance with patient privacy and regulatory standards. After resampling to 1-second and 3-second intervals and filtering out records shorter than one hour, 1,083 records remain. Data collection spans from February 27, 2023, to August 4, 2023. The VitalDB dataset initially includes 6,388 records, resampled at 3-second intervals. After excluding records with more than 20% missing data, 1,522 records remain. To ensure robust model evaluation, both datasets follow an 80%-10%-10% split for training, validation, and testing while maintaining temporal consistency. The CH-OPBP dataset supports multiple prediction horizons, whereas the VitalDB dataset focuses on 3-second sampled data. Table 1 summarizes data partitioning across different sampling rates and prediction scopes. Figure 4 visualizes the partitioning strategy, segmenting each patient’s surgical timeline for structured training and evaluation.
@@
Compared Baselines.
To evaluate HMF, we compare it with traditional and deep learning-based forecasting models. For traditional methods, we use ARIMA Ariyo et al. (2014) and Logistic Regression. ARIMA, a statistical time series model, is trained on 0.5% of the test data for efficiency. Logistic Regression, a standard classifier, is enhanced with 1,566 features extracted via tsfresh [^4] https://github.com/blue-yonder/tsfresh and evaluated on 1% of the dataset. For deep learning baselines, we include LSTM Graves and Graves (2012), Transformer Vaswani et al. (2017), Informer Zhou et al. (2021), and DLinear Zeng et al. (2023). LSTM captures sequential dependencies via recurrence, while Transformer leverages self-attention for long-range interactions. Informer improves efficiency with sparse attention, making it suitable for long-sequence forecasting. DLinear, a lightweight alternative, models trend and seasonal components separately. Further implementation details and hyperparameter settings are in the supplementary material.
Implementation and Evaluation Details.
The HMF model predicts future MAP values based on physiological forecasts of both MAP and SBP, utilizing a 15-minute context window as input. Predictions extend over 5, 10, and 15 minutes to capture short- and long-term trends. Model training minimizes Mean Squared Error (MSE), with evaluation using both MSE and Mean Absolute Error (MAE), focusing on hypotensive segments ($L_{\text{actual}}=1$). Although both MAP and SBP series are predicted, IOH assessment relies solely on the forecasted MAP trajectory. Following clinical guidelines Wesselink et al. (2018), a hypotensive event is identified when predicted MAP remains below $\theta_{\text{MAP}}=65$ mmHg for at least 1 minute. To mitigate the impact of transient fluctuations, a 2-minute skip window Wijnberge et al. (2020) is applied, ensuring evaluation focuses on sustained hypotensive trends rather than isolated drops. Performance is assessed using accuracy, recall, and Area Under the Curve (AUC), with AUC serving as the primary metric for evaluating the model’s ability to differentiate hypotensive trends. Algorithm 1 details the IOH event detection process by comparing predicted and actual MAP sequences in a time-series forecasting framework. Further methodological details, including ground truth labeling and evaluation protocols, are provided in supplement material part.

### v4

@@ Abstract
Intraoperative hypotension (IOH) prediction using past physiological signals is crucial, as IOH may lead to inadequate organ perfusion and significantly elevate the risk of severe complications and mortality. However, current methods often rely on static modeling, overlooking the complex temporal dependencies and the inherently non-stationary nature of physiological signals. We propose a Hybrid Multi-Factor (HMF) network that formulates IOH prediction as a dynamic sequence forecasting task, explicitly capturing both temporal dependencies and physiological non-stationarity. We represent signal dynamics as multivariate time series and decompose them into trend and seasonal components, enabling separate modeling of long-term and periodic variations. Each component is encoded with a patch-based Transformer to balance computational efficiency and feature representation. To address distributional drift from evolving signals, we introduce a symmetric normalization mechanism. Experiments on both public and real-world clinical datasets show that HMF significantly outperforms competitive baselines. We hope HMF offers new insights into IOH prediction and ultimately promotes safer surgical care. Our code is available at https://github.com/Mingyue-Cheng/HMF.
@@ 1 Introduction
Figure 1: Illustration of MAP sequence: temporal dynamics and distribution shifts in intraoperative hypotension events.
@@
Intraoperative mortality has decreased by a factor of $100$ over the past century, making deaths during surgery a rare occurrence Li et al. (2009); Xue et al. (2022). However, mortality within the first month following surgery remains a significant concern, with approximately 2% of patients undergoing inpatient noncardiac surgery dying within $30$ days postoperatively Spence et al. (2019)—amounting to more than $4$ million deaths worldwide each year Saugel and Sessler (2021). These postoperative deaths are most strongly linked to complications Lee et al. (2017), which are often triggered by intraoperative events. Among these, intraoperative hypotension (IOH) is a prevalent and clinically significant complication Kim et al. (2023), characterized by a sustained reduction in mean arterial pressure  [^2] MAP, derived from arterial blood pressure (ABP), is a key indicator of tissue perfusion. It is calculated as $MAP=DBP+\frac{1}{3}(SBP-DBP)$ Gauer (1960), where systolic blood pressure (SBP) and diastolic blood pressure (DBP) represent the maximum and minimum arterial pressures within a cardiac cycle, respectively. (MAP).
@@
Previous studies Hatib et al. (2018a) have shown that IOH events can be predicted using machine learning applied to physiological signals. Clinical trials further indicate that early warnings and timely interventions help reduce hypotension severity and postoperative complications Hwang et al. (2023); Fernandes et al. (2021). Recent advances primarily rely on handcrafted features combined with classifiers such as logistic regression and random forests Lee et al. (2021); Davies et al. (2020), but face two key limitations: (1) limited ability to capture temporal dependencies, and (2) lack of flexibility to handle complex real-world IOH scenarios. Additionally, fixed threshold-based IOH definitions may fail to reflect risk accurately in hypertensive or elderly patients, where more nuanced criteria are often needed.
@@
To overcome these limitations, we reformulate IOH prediction as a multivariate time series forecasting problem. This formulation captures the temporal dynamics and evolving distributions inherent in physiological signals. Since hypotensive events are defined based on MAP series, accurate forecasting directly enhances risk identification. However, as illustrated in Figure 1, several challenges arise. First, MAP waveforms comprise trends and periodic components, which may require separate modeling for improved accuracy. Second, redundancy in long MAP sequences increases computational cost and complicates representation learning. Third, sudden fluctuations alter statistical properties such as mean and standard deviation, making it difficult to maintain consistent performance. Addressing these challenges is essential for a robust and adaptive IOH prediction framework.
@@
We propose the Hybrid Multi-Factor (HMF) framework, which explicitly models temporal structure and mitigates distribution shifts. HMF employs sequence decomposition to separate trend and seasonal components, yielding a more structured representation of historical dynamics. By incorporating multi-factor physiological signals such as MAP and SBP, HMF captures both short-term fluctuations and long-range dependencies. A patch-based Transformer encoder is adopted for efficient sequence representation, while a symmetric normalization mechanism compensates for evolving input distributions. Extensive experiments on public and private intraoperative datasets demonstrate that HMF consistently improves predictive performance and effectively models multi-factor temporal interactions. The key contributions are summarized as follows:
@@
2 Related Work
2.1 Intraoperative Hypotension Forecasting
Existing methods for intraoperative hypotension (IOH) prediction predominantly adopt a static modeling approach, extracting handcrafted features from physiological signals rather than directly modeling their temporal dynamics. Early studies focused on high-fidelity arterial pressure waveforms, leading to the Hypotension Prediction Index (HPI) Hatib et al. (2018b). Machine learning approaches, including ensemble methods Cherifa et al. (2020) and gradient boosting Kendale et al. (2018), integrated preoperative and intraoperative factors but largely treated data as independent samples, overlooking temporal dependencies. Deep learning models, such as RNN-based predictors Jeong et al. (2019) and attention-based frameworks Lu et al. (2023), improved performance by capturing sequential patterns but primarily relied on fixed-length feature representations, failing to fully model non-stationary hemodynamic changes. Recent interpretable models Ritter et al. (2023) enhanced clinical applicability but remained feature-driven. Our work addresses these limitations by formulating IOH prediction as a multivariate time-series forecasting problem, explicitly capturing temporal dependencies and non-stationary dynamics in arterial blood pressure signals.
2.2 Time Series and Sequence Forecasting
Time series forecasting is fundamental in domains such as finance, healthcare, and energy Cheng et al. (2025, 2024). Traditional models like ARIMA Hyndman and Khandakar (2008) and exponential smoothing struggle with high-dimensional physiological signals, while deep learning methods, including LSTMs Graves and Graves (2012) and GRUs Dey and Salem (2017), improve long-term dependency modeling. Recent Transformer-based models enhance long-sequence forecasting effectiveness. Informer Zhou et al. (2021) optimizes self-attention mechanisms through sparse attention and Autoformer Chen et al. (2021) leverages series decomposition to explicitly model trend and seasonal components, which enhances interpretability and improves forecasting stability over long horizons. Lightweight alternatives, such as DLinear Zeng et al. (2023), demonstrate competitive performance by preserving temporal structures with simpler architectures. Due to the data sparsity issue Cheng et al. (2023), GPHT Liu et al. (2024) further introduces generative pretraining to improve forecasting transferability across datasets and horizons. While significant progress has been made in deterministic forecasting, probabilistic approaches such as diffusion-based models Tashiro et al. (2021) Zhang et al. (2025) are also gaining increasing attention for their ability to capture uncertainty. Complementarily, supervised architectures such as TimeDART Wang et al. (2025) highlight the potential of discriminative learning in capturing temporal patterns effectively. On the other hand, the introduction of LLM-based models Jin et al. (2023) Zhou et al. (2023) has enabled the integration of multimodal information. These advancements collectively provide a solid foundation for more accurate modeling of complex medical time series.
While existing models often struggle with real-time adaptability and the challenges posed by non-stationary physiological sequences in dynamic clinical settings, our work bridges this gap by combining efficient patch-based sequence encoding with structured modeling of physiological signals. This design enables rapid inference and robust adaptation to evolving patient states, thereby improving IOH prediction accuracy in dynamic clinical settings.
@@ 2.2 Time Series and Sequence Forecasting
Figure 3: Proposed hybrid multi-factor model for intraoperative hypotension prediction.
@@ 3 Preliminaries
3.1 Multi-factor Sequence Construction
To predict intraoperative hypotension (IOH), we analyze arterial blood pressure (ABP) dynamics to identify early indicators that precede hypotensive events, thereby enabling timely clinical intervention and risk mitigation. During surgery, ABP is continuously monitored at high frequency, capturing critical hemodynamic fluctuations. Prior studies have demonstrated a strong correlation between short-term variations in ABP and the onset of IOH.
Clinically, IOH is defined as a period when the mean arterial pressure (MAP) remains below 65 mmHg for at least one minute. Systolic blood pressure (SBP), the peak value in a cardiac cycle, is also considered due to its physiological relation to MAP. These two signals, MAP and SBP, serve as the primary predictive inputs for IOH forecasting. To construct reliable input sequences we extract clean ABP segments using a sliding window, remove artifacts and compute SBP and diastolic blood pressure (DBP), as shown in the Figure 2.
@@ 3.2 Problem Definition
Building on the multi-factor feature construction framework, we formulate IOH risk prediction as a dynamic sequence forecasting task. Given real-time ABP monitoring, the goal is to forecast future mean arterial pressure (MAP) trends to enable early warnings of hypotensive events. Let $\mathbf{X}_{\text{MAP}}=\{x_{\text{MAP}}^{1},\dots,x_{\text{MAP}}^{T}\}$ and $\mathbf{X}_{\text{SBP}}=\{x_{\text{SBP}}^{1},\dots,x_{\text{SBP}}^{T}\}$ represent historical MAP and systolic blood pressure (SBP) sequences over a context window of length $T$. The objective is to predict MAP values $\mathbf{\hat{Y}}_{\text{MAP}}=\{\hat{y}_{\text{MAP}}^{T+1},\dots,\hat{y}_{\text{MAP}}^{T+\tau}\}$ over horizon $\tau$. A hypotensive event is flagged if predicted MAP drops below a threshold $\theta_{\text{MAP}}$. Further details on annotation and thresholding are provided in the supplementary material. By modeling multi-factor temporal dependencies, our approach improves IOH risk assessment and supports timely clinical decisions.
@@ 4.2 Symmetric Normalization
where $\mu_{X}$ and $\sigma_{X}$ are the mean and standard deviation of $X$. This transformation allows the model to focus on relative changes rather than absolute values, reducing the impact of abrupt fluctuations. Additionally, by normalizing across instances, the method mitigates inter-patient variability, ensuring stable feature distributions.
@@ 4.3 Decomposition Layer
4.4 Representation Layer
@@
To model the intricate dynamics of MAP and SBP series during surgery, we apply a patch embedding technique Nie et al. (2022) after the decomposition layer. This reduces sequence length, lowering computational complexity and improving efficiency. Patch-based modeling captures local patterns while mitigating noise and outliers, enhancing prediction robustness.
@@
$W_{\text{patch}}=\text{Conv1D}(W),$ (5)
where Conv1D applies three consecutive 1D convolutions to $W$. The embedding dimension $d_{\text{model}}$ is chosen to extract key features, improving IOH prediction.
@@
Capturing cross channel dependencies is essential for accurate prediction of multivariate physiological time series Han et al. (2024). Effective IOH prediction requires capturing both short-term and long-term dependencies in MAP and SBP series. The Transformer encoder, with its self-attention mechanism, is well-suited for this task, dynamically weighting the importance of different time segments to model complex temporal patterns.
@@ 4.5 Prediction Layer
To ensure the temporal patterns captured by the Transformer encoder are effectively utilized in MAP predictions, we employ a linear layer to map the sequence representation $Z$ to the predicted output. This transformation allows the model to preserve sequence dependencies while translating extracted features into precise forecasts. To maintain temporal coherence, we use a recurrent forecasting strategy where each predicted step conditions the next, ensuring consistency across the forecasted sequence. The mapping from $Z\in\mathbb{R}^{\frac{L}{S}\times d_{\text{model}}}$ to $Y\in\mathbb{R}^{\frac{L}{S}\times\frac{2TS}{L}}$ is formulated as:
@@ 4.6 Optimization Strategies
To improve prediction accuracy, we integrate a patch-based autoregressive method into our model. This approach begins with the final patch of decomposed component $W$, denoted as $Y_{0}=W_{\frac{L}{S}}$, and sequentially generates subsequent patches using the autoregressive formula:
@@ 5.1 Experimental Setup
We conduct experiments on two real-world clinical datasets: CH-OPBP and VitalDB Lee (2018). The CH-OPBP dataset is collected from real-world intraoperative monitoring records of 3,422 patients, sampled continuously at 100Hz. The data acquisition process strictly adheres to ethical guidelines and protocols approved by the hospital ethics committee, ensuring full compliance with patient privacy regulations and relevant clinical research standards. After resampling to 1-second and 3-second intervals and filtering out records shorter than one hour, a total of 1,083 valid records remain. The data acquisition process also strictly follows patient de-identification protocols and has been formally approved by the Institutional Review Board of Anhui Provincial Hospital. Data collection spans the period from February 27, 2023, to August 4, 2023. The VitalDB dataset initially includes 6,388 intraoperative records, resampled uniformly at 3-second intervals. After excluding records with more than 20% missing data, 1,522 records are retained for use. To ensure reliable model evaluation, both datasets adopt an 80%-10%-10% split for training, validation, and testing, while preserving temporal consistency. The CH-OPBP dataset supports multiple prediction horizons to evaluate forecasting flexibility, whereas the VitalDB dataset focuses specifically on 3-second resolution data. Table 1 summarizes data partitioning across different sampling rates and prediction scopes. Figure 4 visualizes the partitioning strategy, segmenting each patient’s surgical timeline for structured model training and evaluation.
@@
Compared Baselines.
To evaluate HMF, we compare it with traditional and deep learning-based forecasting models. For traditional methods, we use ARIMA Hyndman and Khandakar (2008) and Logistic Regression. ARIMA, a statistical time series model, is trained on 0.5% of the test data for efficiency. Logistic Regression, a standard classifier, is enhanced with 1,566 features extracted via tsfresh [^3] https://github.com/blue-yonder/tsfresh and evaluated on 1% of the dataset. For deep learning baselines, we include LSTM Graves and Graves (2012), Transformer Vaswani et al. (2017), Informer Zhou et al. (2021), and DLinear Zeng et al. (2023). LSTM captures sequential dependencies via recurrence, while Transformer leverages self-attention for long-range interactions. Informer improves efficiency with sparse attention, making it suitable for long-sequence forecasting. DLinear, a lightweight alternative, models trend and seasonal components separately. Further implementation details and hyperparameter settings are in the supplementary material.
Implementation and Evaluation Details.
The HMF model predicts future MAP values based on physiological forecasts of both MAP and SBP, utilizing a 15-minute context window as input. Predictions extend over 5, 10, and 15 minutes to capture short- and long-term trends. Model training minimizes Mean Squared Error (MSE), with evaluation using both MSE and Mean Absolute Error (MAE), focusing on hypotensive segments ($L_{\text{actual}}=1$). Although both MAP and SBP series are predicted, IOH assessment relies solely on the forecasted MAP trajectory. Following clinical guidelines Wesselink et al. (2018), a hypotensive event is identified when predicted MAP remains below $\theta_{\text{MAP}}=65$ mmHg for at least 1 minute. To enable effective early warning and facilitate clinical decision-making, a 2-minute skip window Wijnberge et al. (2020) is applied. Performance is assessed using accuracy, recall, and Area Under the Curve (AUC), with AUC serving as the primary metric for evaluating the model’s ability to differentiate hypotensive trends. Algorithm 1 details the IOH event detection process by comparing predicted and actual MAP sequences in a time-series forecasting framework. Further methodological details, including ground truth labeling and evaluation protocols, are provided in supplement material part.